# thanks: These two authors contributed equallythanks: These two authors contributed equally

### v2

@@ Abstract
Figure 4: Scheme for integrated polarization-based photon-mediated entanglement generation. Because in the typical Bell-state analyzers used, orthogonal polarizations do not interfere, collection of $\hat{\pi}$ and $\hat{\sigma}$-polarized photons into the same modes of separate waveguides, in essentially a dual-rail encoding of the emitted photon state from either atom, allows for the required single-photon interference realized with simply two on-chip 50/50 beamsplitters (BS) acting only on the fundamental quasi-TE waveguide mode.
@@
Our work points to the requirement for optimizing the field output of a collection device projected along the radiating dipole polarization, to optimize photon collection into a particular mode. This offers a simple figure of metric for optimization, including via inverse design lalau2013adjoint , quantification of coupling to undesired modes and polarizations, and of positioning errors on $\eta$. We relate this efficiency to the waists of Gaussian modes, defining spot sizes required to achieve a particular $\eta$. For tight focuses, we show that Gaussian beams perform near (80-90%) optimally and qualitatively resemble the optimal beams. For practically achievable focuses, we predict $\eta$ on the level of multiple percent. However we note that significant enhancements in entanglement rate are not expected for integrated implementation as compared to high-NA bulk optics in state-of-the-art experiments stephenson2020high ; saha2024high . In fact our work points to improvements in design required to achieve collection efficiencies into integrated elements competitive with leading bulk optical demonstrations
@@
Integrated collection as analyzed here with efficient collection optics would enable routing to integrated detectors with high modal selectivity and effective background suppression, as well as the use of compact waveguide-coupled integrated detectors with dimensions independent of the collection aperture. While our discussion has focused on diffractive grating couplers for coupling into SM waveguides, similar functionality may be implemented with devices leveraging metasurface techniques as well huang2023leaky ; zhang2020low . For remote entanglement, the parallelizability of integrated collection offers routes to multiplexed generation on multiple ion pairs in general, as well as alternatives to transport-based multiplexing for increased rate given collection efficiency constraints you2024temporally . The robust polarization encoding enabled and ability to utilize single modes of separate collection waveguides is expected to translate into significantly higher robustness and perhaps fidelity. And, collection efficiency associated with a given optic can be doubled by coupling into two symmetrically placed elements, and combining their outputs via a 50/50 waveguide splitter. Additionally, collection into field profiles with zero gradient at the ion location, e.g. into standing wave profiles associated with such symmetrically placed emitters, may in the future further enable suppression of collection of sideband scattered photons mundt2002coupling ; vasquez2023control , which together with the precise polarization and phase control in integrated settings may in the long run assist in optimizing remote entanglement fidelity.
@@ Appendix B Numerical Details
The overlap integral is numerically calculated along a sample grid. We sample the field along the grating plane over a $200$ $\mu$m $\times$ $200$ $\mu$m domain to capture even the widest Gaussian beams we test, with resolution on this grid is 3501 by 3501 points, to be well over the Nyquist sampling rate for our wavelength. We take the geometry of figure 1, with an ion height of $40$ $\mu$m, though we verify numerically that $\eta$ is unaffected by the choice of integration plane, as indicated by equation 4.

### v3

@@ Abstract
Figure 4: Scheme for integrated polarization-based photon-mediated entanglement generation. Because in the typical Bell-state analyzers used, orthogonal polarizations do not interfere, collection of $\pi$- and $\sigma$-polarized photons into the same modes of separate waveguides, in essentially a dual-rail encoding of the emitted photon state from either atom, allows for the required single-photon interference realized with simply two on-chip 50/50 beamsplitters (BS) acting only on the fundamental quasi-TE waveguide mode. Single-photon detectors 1-4 depicted at the right may be integrated or fiber-coupled off-chip as shown.
@@
Our work points to the requirement for optimizing the field output of a collection device projected along the radiating dipole polarization, to optimize photon collection into a particular mode. This offers a simple figure of merit for optimization, including via inverse design lalau2013adjoint , quantification of coupling to undesired modes and polarizations, and of positioning errors on $\eta$. We relate this efficiency to the waists of Gaussian modes, defining spot sizes required to achieve a particular $\eta$. For tight focuses, we show that Gaussian beams perform near (80-90%) optimally and qualitatively resemble the optimal beams. For practically achievable focuses, we predict $\eta$ on the level of multiple percent. However we note that significant enhancements in entanglement rate are not expected for integrated implementation as compared to high-NA bulk optics in state-of-the-art experiments stephenson2020high ; saha2024high . In fact our work points to improvements in design required to achieve collection efficiencies into integrated elements competitive with leading bulk optical demonstrations
@@
Integrated collection as analyzed here with efficient collection optics would enable routing to integrated detectors with high modal selectivity and effective background suppression, as well as the use of compact waveguide-coupled integrated detectors with dimensions independent of the collection aperture. While our discussion has focused on diffractive grating couplers for coupling into SM waveguides, similar functionality may be implemented with devices leveraging metasurface techniques as well huang2023leaky ; zhang2020low ; ollanik2024integrated . For remote entanglement, the parallelizability of integrated collection offers routes to multiplexed generation on multiple ion pairs in general, as well as alternatives to transport-based multiplexing for increased rate given collection efficiency constraints you2024temporally . The robust polarization encoding enabled and ability to utilize single modes of separate collection waveguides is expected to translate into significantly higher robustness and perhaps fidelity. And, collection efficiency associated with a given optic can be doubled by coupling into two symmetrically placed elements, and combining their outputs via a 50/50 waveguide splitter. Additionally, collection into field profiles with zero gradient at the ion location, e.g. into standing wave profiles associated with such symmetrically placed emitters, may in the future further enable suppression of collection of sideband scattered photons mundt2002coupling ; vasquez2023control , which together with the precise polarization and phase control in integrated settings may in the long run assist in optimizing remote entanglement fidelity.
@@ Acknowledgements.
Disclosures: KKM serves as an advisor to Oxford Ionics (I,E).
Disclosures: The authors declare no conflicts of interest.
@@ Appendix B Numerical Details
The overlap integral is numerically calculated along a sample grid. We sample the field along the grating plane over a $200$ $\mu$m $\times$ $200$ $\mu$m domain to capture even the widest Gaussian beams we test, with a resolution of 3501 by 3501 points, to be well over the Nyquist sampling rate for our wavelength. We take the geometry of Fig. 1 with an ion height of $40$ $\mu$m, though we verify numerically that $\eta$ as calculated vectorially is unaffected by the choice of integration plane, as indicated by equation 4.